2. Trevor earns a base monthly salary of $1500 and a 10% commission on all sales, s. Write an equation that represents Trevor's monthly income, I?


1
Expert's answer
2021-04-14T07:37:51-0400

Let the sales for the month be x

commission will be

"=\frac{10}{100}\times x=0.1x"

income =salary + commission

"I=1500+0.1x"